OpenStudy (anonymous):

What is the value of the function y=2x+3 when x = –1? a) 5 b)-5 c)2 d)1

OpenStudy (igreen):

Plug in -1 for 'x': y = 2x + 3 y = 2(-1) + 3 y = -2 + 3 Can you add -2 + 3?

OpenStudy (anonymous):

thx

OpenStudy (igreen):

No problem, that will be your answer..